Solve the inequality

3x-5/x+2 ≥ x-3




1
Expert's answer
2021-09-02T00:28:36-0400

"3x-\frac{5}x+2 ≥ x-3"


Determine the defined range


"3x-\frac{5}x+2 ≥ x-3, x≠0"


Move the expression to the left hand side and change its sign


"3x-\frac{5}x+2 ≥ x-3"


Calculate


"3x-\frac{5}x+5-x ≥0"


Write all numerators above the common denominator


"\frac{3x^2-{5}+5x-x^2 }{x}≥ 0"


"\frac{2x^2-{5}+5x }{x}≥ 0"


Separate into possible cases

"2x^2-5+5x\u22650"

"x>0"


"2x^2-5+5x\u22640"

"x<0"


Solve the inequality for x


"x\in(-∞, \frac{-5-\sqrt{65}}{4})
 ∪
 (\frac{-5+\sqrt{65}}{4},+∞)"

"x>0"



"x\in(\frac{-5-\sqrt{65}}{4}
 
 ,\frac{-5+\sqrt{65}}{4})"

"x<0"


"x\in(\frac{-5+\sqrt{65}}{4},
+∞)"


"x\in(\frac{-5+\sqrt{65}}{4},
0)"


"x\in(\frac{-5-\sqrt{65}}{4},
0)
∪
(\frac{-5+\sqrt{65}}{4},
+∞),x≠0"


Find the intersection of the solution and the defined range


"x\in(\frac{-5-\sqrt{65}}{4},
0)
∪
(\frac{-5+\sqrt{65}}{4},
+∞)"